OpenStudy (anonymous):

If cos t = 3/5 and 3π/2 < t < 2π, compute sin t and cot t.

OpenStudy (anonymous):

sin t=-4/5 cot t=-4/3

OpenStudy (anonymous):

oops...cot t=3/-4 Thanks!! :-)

OpenStudy (anonymous):

Yeap cot t=3/-4 sorry! What the hell was I thinking? Sorry again

OpenStudy (anonymous):

no problem! :-)